OpenStudy (waheguru):

how 2 sovel this (4y^5)^4 -------- 4y^6*4y^7 how to simplify plz help me

OpenStudy (anonymous):

(4y^5)^4 4^4 y^5*4 4^4 y^20 ---------- = ----------- = ---------- = 4^4-2 y^20-13 = 4^2 y^7 4y^6*4y^7 4^2 y^6+7 4^2 y^13

OpenStudy (waheguru):

i got 16y^17????

OpenStudy (anonymous):

final answer cud also be written as 16y^7

OpenStudy (anonymous):

Look at my working........

OpenStudy (waheguru):

ook thx

OpenStudy (anonymous):

I am frm Delhi, India where r u from??